# Post-hoc Spurious Correlation Neutralization with Single-Weight Fictitious Class Unlearning

### v1

@@ 4.2 Neutralizing a Specific Class in Classifier
Figure 5: Orthogonalization of the hyperplane classifying a class w.r.t. the most significant connection’s weight, i.e., $w_{1}$ the only weight. It results in canceling the model’s discriminatory ability.
Recall that spurious features are viewed as (sub-)classes in our method. Hence, we are interested in precisely neutralizing specific classes in a classifier without destroying the classification hyperplanes for the others. To that end, we leverage the SCA-score to select the most important connection that contributes to classifying a specific class as exclusively as possible. We then make the classification hyperplane orthogonal to its initial state w.r.t. the axis corresponding to that connection. Figure 5 illustrates the notion of orthogonalization for the simplest case where there is only one connection, i.e., only one weight involved; and orthogonalizing the hyperplane w.r.t. that negates the discriminative ability of the classifier. We leverage this concept in n-dimensional space in Theorem 2.
@@
Let $y_{i}=\displaystyle{\bm{w}}_{i}^{\intercal}\displaystyle{\bm{x}}+\displaystyle b_{i}$ define the decision hyperplane for class $i$, and $w_{ji}\in\displaystyle{\bm{w}}_{i}$ be the connection weight with significant impact on the classification outcome. The discriminatory effect of $w_{ji}$ can be nullified by applying the following transformation to it:
@@
In the supplementary material, we provide the formal proof of Theorem 2, where we define normal vectors for the original hyperplane for $y_{i}$ and the hyperplane resulting from applying $f(w_{ji})$ in Eq. (7) to the weight $w_{ji}$ while keeping all other weights unchanged. We show that the dot product of the normal vectors for the two planes is zero, i.e., the underlying hyperplanes are orthogonal, while the difference between the normal vectors only contains non-zero coefficient for the axis corresponding the identified connection. In Algorithm 1, we summarize the proposed process for neutralizing a specific class in the classifier.
@@
2:     Compute accumulative activations $\mathcal{A}_{ac.}^{c}(e^{(L)})$ using Eq. 3 for $\forall c\in\mathcal{C}$.
3:     Compute accumulative gradients magnitudes $\mathcal{G}_{ac.}^{c}(e^{(L)})$ using Eq. 4 for $\forall c\in\mathcal{C}$.
4:     Compute CA score $\Gamma_{CA}(e^{(L)})$ using Eq. 5.
5:     Compute SCA score $\Gamma_{SCA}^{c_{r}}(e^{(L)})$ using Eq. 6.
@@
8:  Orthogonalize the hyperplane associated with $c_{r}$ w.r.t. $e^{*}$ following Eq. 7.
4.3 Removing Fictitious Classes
In Algorithm 2, to remove a fictitious class from the real classes of a model $\mathcal{M}$, we create model $\mathcal{M^{\prime}}$ that is a copy of $\mathcal{M}$. We alter the penultimate layer of the copy such that it has the spurious feature as one of its predicted classes. $\mathcal{M^{\prime}}$ remains frozen, except for the weights of its penultimate layer. We fine-tune those weights on the dataset that contains the spurious feature and its corresponding label. The purpose of this fine-tuning is to bottleneck the spurious correlation learned by $\mathcal{M}$ to an identifiable connection in $\mathcal{M^{\prime}}$ our experiments in Sec. 5 show this can be achieved efficiently. Now that $\mathcal{M^{\prime}}$ recognizes the spurious feature as a class, we can remove it using the method in Sec. 4.2.
Since $\mathcal{M^{\prime}}$ is the same as $\mathcal{M}$ until the last layer, the high-level features extracted by both the models are the same. We can apply the same class removal on $\mathcal{M}$ not to remove the complete class, but to remove the impact of the corresponding high-level features. This process can also be interpreted as removing a sub-class from superclasses or removing a fictitious class from the real classes. Either way, it is notable that the method remains post-hoc because it does not require model retraining. The penultimate layer fine-tuning of $\mathcal{M^{\prime}}$ is for weight identification purpose. The weight gets edited using Eq. (7). It is emphasized that we intentionally present Algorithm 2 such that a copy of $\mathcal{M}$ gets created. This is to clearly explain the underlying idea. Directly removing and replacing $\mathcal{M}$’s penultimate layer is a more memory efficient alternate to implement the same concept.
In general, removing a sub-class from the main classes is more challenging than directly removing a main class, as sub-class features might have much more inter-(sub-)class overlap. This is the reason that we continually sought a connection that is not only significant in classifying a certain fictitious class, but also does it as exclusively as possible. Nonetheless, even if we are able to find such a connection, it remains possible that the same connection also contributes to classifying other sub-classes to some extent. To handle that, we define Partial Feature Neutralization (PFN). Conceptually, PFN enables the hyperplane classifying a class to make a controllable arbitrary tilt w.r.t. its initial state. Equation 8 shows the weight update required to comply to PFN for feature neutralization to an arbitrary extent.
@@ 4.3 Removing Fictitious Classes
where $r\in[0,1]$ is the neutralization rate. Using PFN, we can controllably neutralize features to a level that eliminates model’s over-reliance on them, without significantly altering model’s overall performance. This is established quantitatively in our experiments in Sec. 5.3.
@@ 5 Experiments
To evaluate, we first perform experiments on our underlying unlearning method to verify its ability in effective and precise class removal. This is followed by its application in neutralizing spurious correlations. We start by testing the SCA score in Sec. 5.1. In Sec. 5.2, we perform experiments on the more challenging problem of removing subclasses classes from classes, and eventually our proposed method, denoted as Fictitious Class Removal (FCR), is applied to neutralize spurious correlations in Sec. 5.3. Analysis of CA score effectiveness and ablation study is provided in the supplementary material.
For our experiments on (sub-)class removal we define a simple two-layer convolutional model Conv-2Net, and also use ResNet-18 (Sec. 5.1 and Sec. 5.2). Aligned with the literature, ResNet-50 is used for experiments on spurious correlation neutralization (Sec. 5.3).
@@
Figure 7: Average accuracy change after removing randomly classes. Left: CIFAR-10 results using a well-trained Conv-2Net. Right: CIFAR-100 results using ResNet-18. The accuracy for non-removed classes improves as more classes are removed. The accuracy for the removed classes approaches random guessing.
5.1 Class Removal
We use the SCA-score in Eq. 6 to rank weights according to their significance in class-specific classification and apply the transformation in Eq. 7 to unlearn each class separately, without compromising the accuracy of the others.
Figure 6 provides the accuracy curves for editing up to 3 weights from the model for removing CIFAR-10 classes individually. As apparent, class removal for every class dropped the target class accuracy to almost zero with single weight editing. Moreover, removing the class did not negatively impact the accuracies for other classes. For instance, after removing class 0, the accuracy for classes 1, 3, 5, 6, and 7 remained intact, while there is a slight increase in the accuracy for classes 2, 4, 8, and 9.
We also apply our class removal technique to remove multiples classes. Figure 7 summarizes the result for removing up to 8 random classes for CIFAR-10 and 98 classes for CIFAR-100 by editing just one weight. The average accuracy for the removed classes clearly drops to around random-guessing while there is no damage to the un-removed classes. These results clearly establish our approach as an effective class removal strategy.
5.2 Fictitious Class Removal
Following the order of our discussion in Sec. 4, here we examine the notion of removing fictitious classes from real classes. To this end, we use ParityMNIST [26] dataset consisting of two classes; namely, ‘Class 0’ and ‘Class 1’ that respectively contain even and odd digits from 0 to 9. We assume that classes 0 and 1 are the true classes, i.e., superclasses, while the digits within them are the fictitious classes, i.e., subclasses. Following our Algorithm 2, a model $\mathcal{M}$ on ParityMNIST is trained. Then we tune a copy of $\mathcal{M}$ on MNIST. Since the digits are fictitious classes here, we can treat this model as $\mathcal{M^{\prime}}$. We use the information of activations and gradients of the 10 digit classes to compute the SCA-Score for each class. To remove a fictitious class, we perform the class removal as proposed in Sec. 5.1, i.e., using the SCA-score obtained from $\mathcal{M^{\prime}}$ and used for the model $\mathcal{M}$ trained on ParityMNIST.
@@ 5.2 Fictitious Class Removal
Figure 8: Accuracy of removed and retained subclasses on ParityMNIST with different neutralization rate $r$. Left: Only accumulative gradients and activations are used for selecting the most significant connections. Right: SCA-score is used for selecting the connections. Further elaborated in supplementary.
@@ 5.3 Removing Spurious Features
Figure 9: Feature neutralization by different neutralization rate values ($r$). The best $r$ is obtained where the gap between average and worst-group accuracies is minimum, and is shown with a vertical dotted lines for Waterbirds (left), CelebA (middle), and MetaShift (right).
Table 1: Performance comparison of our FCR technique with the state-of-the-art methods, presenting worst-group (Worst) and gap (Gap) between Worst and average accuracies on Waterbirds, CelebA and MetaShift datasets. The best results are shown in bold. Ours is a unique post-hoc method, which uses a single weight to remove spurious correlation for a given feature.
{tblr} column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, cell11 = r=2, cell12 = r=2, cell13 = c=2, cell15 = r=2, cell16 = c=2, cell19 = c=2, cell112 = c=2c, cell311 = c, cell312 = c, cell313 = c, cell42 = c, cell411 = c, cell412 = c, cell413 = c, cell52 = c, cell511 = c, cell512 = c, cell513 = c, cell62 = c, cell611 = c, cell612 = c, cell613 = c, cell72 = c, cell711 = c, cell712 = c, cell713 = c, cell82 = c, cell811 = c, cell812 = c, cell813 = c, cell92 = c, cell911 = c, cell912 = c, cell913 = c, cell102 = c, cell1011 = c, cell1012 = c, cell1013 = c, cell112 = c, cell1111 = c, cell1112 = c, cell1113 = c, cell1211 = c, cell1212 = c, cell1213 = c, cell132 = c, cell1311 = c, cell1312 = c, cell1313 = c, cell142 = c, cell1411 = c, cell1412 = c, cell1413 = c, hline1,15 = -0.08em, hline2 = 3-4,6-7,9-10,12-130.03em, hline3,5,9 = -0.05em, Method & Group Info Post-hoc Waterbirds CelebA MetaShift (Cat vs. Dog)
[…]
Group DRO (ICLR ’20) [32] Yes Yes ✗ 91.4 $\pm$1.1 2.1 88.9 $\pm$2.3 4.0 66.0 $\pm$3.8 7.6
PDE (ICML ’23) [9] Yes Yes ✗  90.3$\pm$0.3 2.1   91.0$\pm$0.4 1.0 - -
JTT (ICML ’21) [22] No Yes ✗ 86.7$\pm$ $N/A$ 6.6 88.0$\pm$ $N/A$ 6.9 64.6 $\pm$2.3 9.8
DFR (ICLR ’23) [17] No Yes ✗ 92.9$\pm$0.2 1.3 88.3$\pm$1.1 3.0 72.8 $\pm$3.8 4.7
LBC (IJCAI ’24) [47] No Yes ✗ 88.1$\pm$1.4 6.0 87.4$\pm$1.8 5.0 - -
DaC (CVPR ’24) [29] No Yes ✗ 92.3$\pm$ 0.4 3.0 81.9 $\pm$ 0.7 9.5 78.3 $\pm$ 1.6 1.0
Base (ERM) No No ✗ 75.3$\pm$ 0.6 24.4 48.8 $\pm$ 1.1 47.1 62.1 $\pm$ 4.8 10.8
MaskTune (NeurIPS ’22) [1] No No ✗ 86.4$\pm$ 1.9 6.6 78.0$\pm$ 1.2 13.3 66.3 $\pm$ 6.3 6.8
LC (ICLR ’23) [23] No No ✗ 90.5$\pm$ 1.1 $N/A$ 88.1$\pm$ 0.8 $N/A$
DISC (ICML ’23) [38] No No ✗ 88.7$\pm$ 0.4 5.1 - - 73.5$\pm$ 1.4 2.0
DFR+ExMap (CVPR ’24) [4] No No ✗ 92.5$\pm$ $N/A$ 3.5 84.4$\pm$ $N/A$ 7.4 - -
FCR - Ours No No ✓ 93.2 $\pm$ 0.3 1.9 84.9$\pm$ 0.7 5.4 78.3 $\pm$ 0.4 0.1
@@
Results: We compare our method with the state-of-the-art methods of this direction, namely Group DRO [32], PDE [9], JTT [22], DFR [17], LBC [47], DaC [29], MaskTune [1], LC [23], DISC [38], and DFR+ExMAP [4]. The comparison results are given in Tab. 1. Our method achieves the highest worst-group result for the Waterbirds and MetaShift datasets among all categories, as specified by the use of Group Information requirement by the methods. On CelebA, our method has a low gap between average and worst-group accuracies. As indicated, our method is the only technique that can be applied post-hoc and it manipulates only one weight per class. Hence, it can be applied to already well-optimized models without requiring updates for excessive number of weights.
@@
Figure 9 comprehensively presents the worst group and average accuracies using different neutralization rates $r$ for all the three datasets. We choose $r$ when the gap between average and worst group accuracies is the least and consider it as the most fair point. As can be seen in Fig. 9, the lowest gap between the accuracies for Waterbirds occurs at $r=13$ (left) and the lowest gap for CelebA is at $r=65$ (middle). The lowest gap for MetaShift takes place at 0.0199, which is as low as $0.1\%$. Please note the worst group accuracy declines after these points because further neutralization changes the worst group to another group that had relied more on the removed feature. Please refer to supplementary material for more results and visualizations. Our implementation will be made public after acceptance.

### v2

@@ 4.2 Neutralizing a Specific Class in Classifier
Recall that spurious features are viewed as (sub-)classes in our method. Hence, we are interested in precisely neutralizing specific classes in a classifier without destroying the classification hyperplanes for the others. To that end, we leverage the SCA-score to select the most important connection that contributes to classifying a specific class as exclusively as possible. We then make the (near) optimal classification hyperplane orthogonal to its initial state w.r.t. the axis corresponding to that connection. Fig. 5 illustrates the notion of orthogonalization for the simplest case where there is only one connection, i.e., only one weight involved; and orthogonalizing the hyperplane w.r.t. that negates the discriminative ability of the classifier. We leverage this concept in n-dimensional space in Theorem 2.
@@
Let $y_{i}=\bm{w}_{i}^{\intercal}\displaystyle\bm{x}+\displaystyle b_{i}$ define the decision hyperplane for class $i$, and $w_{ji}\in\displaystyle\bm{w}_{i}$ be the connection weight with significant impact on the classification outcome. The discriminatory effect of $w_{ji}$ can be nullified by applying the following transformation to it:
@@
Proof.
Let $\bm{u}=[w_{1i},\dots,w_{ni},-1]$ be the normal vector of $y_{i}$ and $\bm{u}^{\prime}$ be the normal vector of $y^{\prime}_{i}$ which is the resulting hyperplane by applying Eq. (16) on $w_{ji}$. The dot product of $\bm{u}$ and $\bm{u}^{\prime}$ is
$\bm{u}\cdot\bm{u}^{\prime}=\sum_{k=1}^{n}w_{ki}^{2}-w_{ji}^{2}+1+(-\frac{\|\bm{w}_{i}\|_{2}^{2}-w_{ji}^{2}+1}{w_{ji}}.w_{ji})=0\quad$ (17)
Hence, the resulting hyperplane is orthogonal to its original state w.r.t. $w_{ij}$. For a well-trained classifier, this implies that the classifier can no longer discriminate between the samples along the $w_{ij}$ dimension after the transformation. Next, we calculate the difference caused by the transformation between the normal vectors of the two hyperplanes:
$\bm{u}-\bm{u}^{\prime}=\bm{v}_{j}(-\frac{\|\bm{w}_{i}\|_{2}^{2}-w_{ji}^{2}+1}{w_{ji}}-w_{ji}),$ (18)
where $\bm{v}_{j}$ is the one-hot vector associated with $x_{j}$ axis. From Eq. (17) and Eq. (18) we conclude that the transformation in Eq. (16) transforms the hyperplane to be orthogonal to its initial state merely with changes occurring along $x_{j}$ axis. ∎
4.3 Removing Fictitious Classes
In Algorithm 2, to remove a fictitious class from the real classes of a model $\mathcal{M}$, we create model $\mathcal{M^{\prime}}$ that is a copy of $\mathcal{M}$. We alter the penultimate layer of the copy such that it has the spurious feature as one of its predicted classes. $\mathcal{M^{\prime}}$ remains frozen, except for the weights of its penultimate layer. We fine-tune those weights on the dataset that contains the spurious feature and its corresponding label. The purpose of this fine-tuning is to bottleneck the spurious correlation learned by $\mathcal{M}$ to an identifiable connection in $\mathcal{M^{\prime}}$ - our experiments in Sec. 4.3 show this can be achieved efficiently. Now that $\mathcal{M^{\prime}}$ recognizes the spurious feature as a class, we can remove it using the method in Sec. 4.2.
Since $\mathcal{M^{\prime}}$ is the same as $\mathcal{M}$ until the last layer, the high-level features extracted by both models are the same. We can apply the same class removal on $\mathcal{M}$ not to remove the complete class, but to remove the impact of the corresponding high-level features. This process can also be interpreted as removing a sub-class from superclasses or removing a fictitious class from the real classes. Either way, it is notable that the method remains post-hoc because it does not require model retraining. The penultimate layer fine-tuning of $\mathcal{M^{\prime}}$ is for weight identification purpose. The weight gets edited using Eq. (16). It is emphasized that we intentionally present Algorithm 2 such that a copy of $\mathcal{M}$ gets created. This is to clearly explain the underlying idea. Directly removing and replacing $\mathcal{M}$’s penultimate layer is a more memory-efficient alternative to implement the same concept.
@@ 4.3 Removing Fictitious Classes
2:  Compute accumulative activations $\mathcal{A}_{ac.}^{c}(e^{(L)})$ using Eq.(12) for $\forall c\in\mathcal{C}$.
3:  Compute accumulative gradients magnitudes $\mathcal{G}_{ac.}^{c}(e^{(L)})$ using Eq. (13) for $\forall c\in\mathcal{C}$.
4:  Compute CA score $\Gamma_{CA}(e^{(L)})$ using Eq. (14).
5:  Compute SCA score $\Gamma_{SCA}^{c_{r}}(e^{(L)})$ using Eq. (15).
@@
8: Orthogonalize the hyperplane associated with $c_{r}$ w.r.t. $e^{*}$ following Eq. (16).
In general, removing a sub-class from the main classes is more challenging than directly removing a main class, as sub-class features might have much more inter-(sub-)class overlap. This is the reason that we continually sought a connection that is not only significant in classifying a certain fictitious class, but also does it as exclusively as possible. Nonetheless, even if we are able to find such a connection, it remains possible that the same connection also contributes to classifying other sub-classes to some extent.
To handle that, we define Partial Feature Neutralization (PFN). Conceptually, PFN enables the hyperplane classifying a class to make a controllable, arbitrary tilt w.r.t. its initial state. Eq. (19) shows the weight update required to comply with PFN for feature neutralization to an arbitrary extent.
@@
where $r\in[0,1]$ is the neutralization rate. Using PFN, we can controllably neutralize features to a level that eliminates model’s over-reliance on them, without significantly altering model’s overall performance. This is established quantitatively in our experiments in 5.3. Optimal $r$ can be simply found using binary search, where we consider an $r$ as excessive if it leads to a high reduction in average performance (compared to the improvement for the worst group) or if it changes the worst group, i.e., the accuracy of the initial worst group overtakes the average or another group’s accuracy.
@@ 5 Experiments
We first evaluate our unlearning method and verify its ability in effective and precise class removal. This is followed by its application in neutralizing spurious correlations. We start by testing the SCA score in Sec. 5.1. In Sec. 5.2, we perform experiments on the more challenging problem of removing subclasses from classes. Finally, our proposed method, denoted as Fictitious Class Removal (FCR), is applied to neutralize spurious correlations in Sec. 5.3. Analysis of CA score effectiveness and ablation study is provided in Sec. 6.2.
For our experiments on (sub-)class removal, we define a simple two-layer convolutional model Conv-2Net, and also use ResNet-18 (5.1 and 5.2). Aligned with the literature, ResNet-50 is used for extensive experiments on spurious correlation neutralization (5.3).
5.1 Class Removal
We use the SCA-score in Eq. (15) to rank weights according to their significance in class-specific classification and apply the transformation in Eq. (16) to unlearn each class separately, without compromising the accuracy of the others.
Fig. 6 provides the accuracy curves for editing up to three weights from the model for removing CIFAR-10 classes individually.
We can see that our class removal method decreases the target class accuracy to almost zero with single weight editing.
Moreover, removing the class does not negatively impact the accuracies for other classes. For instance, after removing class 0, the accuracy for classes 1, 3, 5, 6, and 7 remained intact, while there is a slight increase in the accuracy for classes 2, 4, 8, and 9.
We also apply our class removal technique to remove multiple classes. Fig. 7 summarizes the result for removing up to 8 random classes for CIFAR-10 and 98 classes for CIFAR-100 by editing just one weight per class. The average accuracy for the removed classes clearly drops to around random-guessing while there is no damage to the un-removed classes. These results clearly establish our approach as an effective class removal strategy.
5.2 Fictitious Class Removal
@@
Figure 7: Average accuracy change after removing random classes. Left: CIFAR-10 results using a well-trained Conv-2Net. Right: CIFAR-100 results using ResNet-18. The accuracy for non-removed classes improves as more classes are removed.
Following the order of our discussion in Sec. 4, here we examine the notion of removing fictitious classes from real classes. To this end, we use ParityMNIST [50] dataset consisting of two classes; namely, ‘Class 0’ and ‘Class 1’ that respectively contain even and odd digits from 0 to 9. We assume that classes 0 and 1 are the true classes, i.e., superclasses, while the digits within them are the fictitious classes, i.e., subclasses. Following our Algorithm 2, a model $\mathcal{M}$ on ParityMNIST is trained. Then we tune a copy of $\mathcal{M}$ on MNIST. Since the digits are fictitious classes here, we can treat this model as $\mathcal{M^{\prime}}$. We use the information of activations and gradients of the 10 digit classes to compute the SCA-Score for each class. To remove a fictitious class, we perform the class removal as proposed in Sec.5.1, i.e., using the SCA-score obtained from $\mathcal{M^{\prime}}$ for the model $\mathcal{M}$ trained on ParityMNIST.
@@ 5.3 Removing Spurious Features
Figure 8: Accuracy of removed and retained subclasses on ParityMNIST with different neutralization rate $r$. Left: Only accumulative gradients and activations are used for selecting the most significant connections. Right: SCA-score is used for selecting the connections.
Figure 9: Feature neutralization by different neutralization rate values ($r$). The best $r$ can be found using binary search, where $r$ is considered excessiveexcessive if it reduces the average performance immensely or if it leads to change of the worst group, and is shown with a vertical dotted lines for Waterbirds (left), CelebA (middle), and MetaShift (right).
Figure 10: GradCAM saliency maps for representative samples that are misclassified by ERM model and correctly classified by our method. The visualizations confirm that our method effectively remove the model’s attention from non-robust features.
@@
Results: We compare our method with the state-of-the-art methods of this direction, namely Group DRO [24], PDE [17], JTT [15], DFR [22], LBC [53], DaC [54], MaskTune [16], LC [3], DISC [55], and DFR+ExMAP [33]. The comparison results are given in Table I. Our method achieves the highest worst-group result for the Waterbirds and MetaShift datasets among all categories, whether or not they use Group Annotations, and on CelebA, it has the lowest gap between average and worst-group accuracies in the last category. As indicated, ours is the only method that can be applied post-hoc and it manipulates only one weight per class. Hence, it can be applied to already well-optimized models without requiring updates for excessive number of weights.
@@
Fig. 9 comprehensively presents the worst group and average accuracies using different neutralization rates $r$ for all the three datasets. We can choose optimal $r$ value using binary search as discussed in Sec. 4.3. As observed in 9, the optimal value for Waterbirds is $r=13$ (left) and for CelebA is $r=65$ (middle). The best $r$ for MetaShift is 0.0199 (right), on which the gap between the worst group and average accuracies is as low as $0.1\%$. The decline in worst group accuracy is observed after these points because further neutralization changes the worst group to another group since it still relies more on the removed feature.
To confirm that our technique is able to take away model’s attention from spurious features and re-focus it on the causal features, we also observe the saliency maps of the models before and after our modifications.
In Fig. 10, we present representative qualitative results for Waterbirds, CelebA and MetaShift datasets. From the figure, it is clear that our method is able to fix the spurious correlation based associations in ERM based models very effectively. Since our method eventually disrupts the weight in the penultimate layer, it remains largely agnostic to backbone architectures. Our implementation will be made public after publication.
TABLE I: Performance comparison of our FCR technique with the state-of-the-art methods, presenting worst-group (Worst) and gap (Gap) between Worst and average accuracies on Waterbirds, CelebA and MetaShift datasets. The best results are shown in bold. Ours is a unique post-hoc method, which uses a single weight to remove spurious correlation for a given feature.
[…]
TABLE II: Details of fine-tuning $\mathcal{M}^{\prime}$ for different dataset.
Waterbirds CelebA MetaShift # of epochs 5 5 100 # of params updated 4048 4048 4048 Percentage of params updated $\sim 1.7\times 10^{-2}\%$ $\sim 1.7\times 10^{-2}\%$ $\sim 1.7\times 10^{-2}\%$ Number of training samples 16,788 202,599 749 Learning rate $5\times 10^{-6}$ $10^{-3}$ $10^{-3}$ Optimizer Adam Adam Adam Batch size 32 32 16 Accuracy on new task $92.85\%$ $84.40\%$ $84.88\%$
6 Further Discussions
6.1 Experimental Setting
In Sec. 4.3, we talked about the copy $\mathcal{M}^{\prime}$ of the original model $\mathcal{M}$. In Algorithm 2, we also noted ‘fine-tuning’ of $\mathcal{M}^{\prime}$. Here, we give further details about the fine tuning, as summarized in Table II. The table summarizes the hyper-parameter settings used for fine-tuning model $\mathcal{M}^{\prime}$ for the three datasets used. The training data used for fine-tuning for CelebA is produced simply by using gender targets. For Waterbirds we used the ‘Water’ and ‘Land’ background images from Places365 [56] and used the foreground segmentation masks provided by [24]. For MetaShift, we utilized the segmentation technique proposed in [57] for generating foregrounds, and then filled the remaining image using [58]. A few representative samples of the produced foreground-background dataset are shown in Fig. 11.
Figure 11: Representative samples of the prepared foreground-background dataset to fine-tune $\mathcal{M}^{\prime}$ for MetaShift.
6.2 Ablation Study on CA-Score
In Sec. 4.1, we introduced CA-score. This score is compositional in its nature, comprising components related to activations and gradients. We test the importance of the weights selected by the individual components of the CA-score and the eventual CA-score proposed in the paper.
Figure 12: Ablation for CA-Score. The significance of edges ranked by our composite CA-Score is considerably higher than to any of its constituent terms. The unlearning based on CA-Score is more effective.
Figure 13: Unlearning based on CA-Score compared to other common metrics of computing edge importance. (Left) Results on CIFAR-10. (Right) Results on CIFAR-100.
Fig. 12 gives the comparison results of the CA-Score and its constituent components, namely entropy of activations over classes ($H(\bigcup_{c\in\mathcal{C}}\mathcal{G}_{acc.}^{\mathcal{C}}(e_{ji}^{(l)}))$) and the inverse of entropy of gradients over classes ($\frac{1}{H(\bigcup_{c\in\mathcal{C}}\mathcal{A}_{acc.}^{\mathcal{C}}(e_{ji}^{(l)}))}$) on our Conv-2 model trained on CIFAR-10. It is evident that CA-score, as a composite metric, clearly achieves more favorable results in impacting the model performance through the weights specified by its eventual computation.
We also take a step further and compare the performance of removing weights using our CA-score with popular exiting techniques of using $L_{1}$-norm and Hessian to specify the most significant model weights.
Figure 14: Changes of the worst group on MetaShift. Excessive neutralization rate initially changes the worst group from Dog to Cat, and eventually larger values corrupt the model.
In Fig. 13, results are shown for Conv-2Net on CIFAR-10 dataset and ResNet-18 on CIFAR-100 dataset. We gradually remove the most important weights suggested by our CA-score and the existing techniques, i.e., using $L_{1}$-norm, Hessian and $L_{1}$-norm times Hessian. It is observable that the CA-score generally perform much better than using conventional methods to specify the most important weights.
6.3 Details on Subclass Removal on ParityMNIST
Fig. 8 presented results on ParityMNIST. To further elaborate on these results, we expand on them in Fig. 15. It can be seen that with the exception of sub-class 3, all other subclasses are removed effectively with relatively low compromise in the accuracy for other subclasses.
We conjecture that the reason for low performance (i.e., higher accuracy despite removal) on subclass 3 stems in the fact that this class does not truly have exclusive high-level features as other digits. For example, subclass 3 shares its high-level features extensively with sub-class 8. Hence, the objective of neutralizing 3 exclusively remains ill-formed. This also highlights an important fact that our general approach of editing the model with limited (or just one) weight may face restrictions if the features of the underlying classes are not exclusive to other classes.
Figure 15: Performance of subclass removal on parityMNIST. The dashed red and the solid gray curves represent the removed subclass accuracy and the non-removed subclasses accuracy average, respectively. The subclasses are removed effectively without compromising other subclasses, except for digit ‘3’, which could not be removed completely. We conjecture that this is due to the fact that ‘3’ lacks exclusive high-level features. The set of high-level features of ‘3’ is a subset of the set of the high-level features of ‘8’.
Figure 16: t-SNE visualization of features for Waterbirds, CelebA, and MetaShift. Waterbirds and MetaShift show more disentagled embeddings compared to CelebA, which result in relatively lower optimal $r$ and higher spurious features neutralization performance on these datasets.
6.4 Changes of the Worst Group
Extremely high neutralization rates lead to change in the worst group and eventually corruption of the model. Fig. 14 gives the group-wise accuracy for the case of Metashift with different values of $r$. The average accuracy and the accuracies for each of the classes increase for $r$ values up to approximately 0.01, which shows the proposed method is beyond merely establishing a tradeoff between the accuracies of different groups.
6.5 t-SNE Visualizations
For better understanding of why the optimal neutralization rate ($r$) differs among datasets, and why our method is more successful on some datasets compared to others, we utilized t-SNE visualizations (see Fig. 16). The visualizations show more separable spurious features for Waterbirds and MetaShift in comparison to CelebA in which features are more intertwined. This observation is well aligned with higher $r$ needed for our method for CelebA and better results for Waterbirds and MetaShift, as the spurious features can be neutralized more precisely. Additionally, it shows the potential of well-disentangled embeddings to help the proposed method to mitigate spurious features.
6.6 Independence from Underlying Architectures
In this paper, we conducted experiments on ResNet architecture for compliance with the literature. Nevertheless, we claimed that the proposed spurious correlation is agnostic to the model architecture as it works on the last fully connected layer. Our method tries to neutralize the high-level spurious features that are extracted by the preceding layers. So, it is not affected by different procedures of inference in various architectures. To verify this claim experimentally, we adapted our FCR technique to ViT-B/16 models trained on Waterbirds and CelebA datasets taken from [59] (see Table III).
TABLE III: Results on ViT-B/16. Worst group and average accuracies are reported for Waterbirds and CelebA. The improvement on worst groups while low degradation in averages verifies the effectiveness of our method on ViT architecture.
Waterbirds CelebA
Worst (%) Avg. (%) Worst Avg. (%)
ERM 89.30 $\pm$1.95 96.75 $\pm$0.05 94.10 $\pm$0.51 97.40 $\pm$0.42
Ours 93.55 $\pm$ 0.89 96.16 $\pm$0.16 95.42 $\pm$0.58 97.08 $\pm$ 0.79